OpenStudy (moneydaughter1):

FAN AND MEDAL 8-2b= -2/3(12b+6)

OpenStudy (anonymous):

I assume you are trying to solve for b?

OpenStudy (moneydaughter1):

Yep

OpenStudy (anonymous):

The trick to begin with is to insolate the values of b on one side of the equation and the rest of the numbers on the other side. Use the property of distribution [ a(b+c)=ab+ac ]

OpenStudy (anonymous):

Distribute the-2/3 first. You'll get: 8-2b= (-8b+4) Combined like terms to get the B by itself on one side. 8=(-6b+4) 4=(-6b) Then you can just divide to get B

OpenStudy (moneydaughter1):

We need help on the very last step which is 8-2b=-8b+-4. What do you do after that?

OpenStudy (anonymous):

You will move all of your b values to one side of the equals sign. Then move your number values to the other side. For example: a+4=2a-5 becomes a-2a = -5 -4

OpenStudy (moneydaughter1):

Wouldn't it be -10b because you are adding two negatives?

OpenStudy (moneydaughter1):

@Nnesha Please help me

OpenStudy (moneydaughter1):

Is the answer -1.5?

OpenStudy (anonymous):

So the main goal to begin with is isolating the variable. Whenever you solve for a variable, you need to isolate it. You do so using the subtraction or addition property of equality. Its like this: 5=2+b In order to solve that, you need to isolate the variable. You of course know that it's three, but you solve it by subtracting 2 from the five. It's the same with subtraction, except you add it. If a number has a negative sign, you'll add it to the other side of the equation. If it has a positive sign (or no sign) then it needs to be subtracted from the opposite side. 8-2b= -2/3(12b+6) So you know to distribute the -2/3. So now you have this. 8-2b=-8b-4 ( i made a mistake in my first comment, so you can just ignore that) You need to isolate the variable, so you need to move them all to one side. We can do that by -2b to the -8b. That's why it's not -10b. It's -6b because youre adding the number. Heres an example of what i mean: -5= -2+b You know that it is -3, but if you add them the way you're trying to, you'll get b= -7. Anyways, you have the variable to one side now. 8=-6b-4 Add the 4 to the other side to finally isolate the variable completely. 12= -6b. Now all you have to do is solve for b. You'll do that by dividing by -6, since that is the coefficient to the B. You'll get b=(-2). Remember you can check it by substituting it into the original problem. 8-2(-2)= -2/3(12(-2)+6) Use PEMDAS 8+4=-2/3(24+6) 12=16-4 12=12. I hope this helped clear it up

OpenStudy (moneydaughter1):

Thx. Here is your medal and I am a fan of you.